KyledaGreat:

Solve the following exponential equation. Express your answer as both an exact expression and a decimal approximation rounded to two decimal places. 7^−2x+9 = 608

KyledaGreat:

1 attachment
KyledaGreat:

@vocaloid

Vocaloid:

two things - there's an extra 2 at the end that shouldn't be there I assume you got to this part correctly -2x = ln(608)/ln(7) - 9 dividing both sides by -2 gives us x = 9/2 - ln(608)/ (2 * ln(7)) also, it says to round to 2 decimal places, so 2.8529108 needs to be rounded to 2.85

KyledaGreat:

do it what i have on the left is right ?

Vocaloid:

you need to get rid of that last 2 at the end, then it's right

KyledaGreat:

oh i'm sorry you right

KyledaGreat:

Solve the following exponential equation. Express your answer as a decimal approximation rounded to two decimal places. e^x−7 = 84x x = __________

Vocaloid:

is the x in 84 x also an exponent?

KyledaGreat:

yes

KyledaGreat:

1 attachment
Vocaloid:

e^(x-7) = 8^(4x) taking the natural log of both sides will eliminate the e on the left side and give us ln(8^(4x)) on the right side x - 7 = ln ( 8^(4x) ) using our log rules, you can bring that 4x down from inside the log, to right in front of the log x - 7 = 4x * ln(8) can you take it from here? move your x terms to one side, factor out x, solve for x

KyledaGreat:

x - xln (2097152)

Vocaloid:

hm? how did you get this answer?

KyledaGreat:

by simplifying

Vocaloid:

would recommend reviewing how to solve 1-variable equations when you get a chance from here: x - 7 = 4x * ln(8) we can move our x-terms to one side by subtracting 4x*ln(8) from both sides we can also add 7 to both sides to move the 7 to the other side, giving us: x - 4x(ln8) = 7 factoring out x x(1-4ln(8)) = 7 dividing both sides by (1-4ln(8)) x = 7/(1-4ln(8))

KyledaGreat:

-0.95657606987

Vocaloid:

yeah just remember to round to 2 places

KyledaGreat:

-0.95 ?

Vocaloid:

review rounding when you get a chance -0.956... look at the 2nd decimal place. then we look at the next number, 6. because 6 is greater than 5, we round up to -0.96

KyledaGreat:

oh i see

KyledaGreat:

Solve the following exponential equation. Express your answer as a decimal approximation rounded to two decimal places. e^2x−3 = 13^x/5

KyledaGreat:

x = 2.01747112

Vocaloid:

good, but remember to round to 2 places 2.017 ---> 7 is greater than 5 so we round up to 2.02

KyledaGreat:

1 attachment
KyledaGreat:

just you can see how it looks